OpenStudy (anonymous):

If a = 3 and b = 5, find a/a+b=___?

OpenStudy (anonymous):

3/3 + 5 1 + 5 6

OpenStudy (anonymous):

3/8

OpenStudy (anonymous):

2/8 3/8 3/5

OpenStudy (anonymous):

what about find b/a+b=___?

OpenStudy (anonymous):

same answer?..cus these are the answer choices 8/5 3/8 5/8

OpenStudy (anonymous):

5/8

OpenStudy (anonymous):

what is 7/8 divided 1/2

OpenStudy (anonymous):

6 is the answer. 3/3+5 =1+5 =6

OpenStudy (anonymous):

answer was not 3/5

OpenStudy (anonymous):

It's six?